In how many ways can a committee of three men and four women be formed from a group of 11 men and 11 women?

Answer by ewatrrr(24785)

 
Hi
In how many ways can a committee of three men and four women be formed
from a group of 11 men and 11 women?
11C3*11C4 = 165*330 = 54,540 ways